OpenStudy (1018):

where is the curve x^(2)y - xy - x^(2) +10x - 12y -21 = 0 symmetric?

OpenStudy (amistre64):

what conic section is it?

OpenStudy (amistre64):

hmm, thought it was a conic at first glance

OpenStudy (1018):

i'm sorry im not really sure, the question is just that one.

OpenStudy (1018):

i already forgot this one, sorry

OpenStudy (amistre64):

it is a question from your course right? if so then you must have material that tells you how to approach it

OpenStudy (amistre64):

if it posted correctly, then the graph from the wolf doesnt show it as a conic the x^2y throws it off

OpenStudy (amistre64):

Ax^2 + By^2 +Cx + Dy+ Exy +F = 0 i believe is the form of a rotated conic

OpenStudy (amistre64):

highest degree in this questions is 3

OpenStudy (1018):

there are actually choices. it's x axis, y axis, origin and none

OpenStudy (kainui):

Did you just change the question?

OpenStudy (amistre64):

ahh, then symmetry test is it

OpenStudy (1018):

i apologize but my memory of this math is little. haha that's why im reviewing lol

OpenStudy (1018):

no those are just the choices for the same quesiton

OpenStudy (amistre64):

how do we test those choices?

OpenStudy (astrophysics):

Oh yay I was right, I thought I read the question wrong lol, so symmetry test does work?

OpenStudy (amistre64):

if symmetric about the y axis: then f(-x) = f(x) if symmetric about the x axis: then f(-x) = -f(x) if niether, then none

OpenStudy (amistre64):

i was thinking an axis of symmetry question meself :)

OpenStudy (1018):

so can you teach me what to do? haha